OpenStudy (anonymous):

g(n) = n^2 - 5n h(n) = 2n + 1 Find g(y-2) - h(y-2) I know how to do this problem, but I don't understand why the result is y^2-11y+17. I get y^2-11y+11 First, when you plug in y-2 into g(y-2)=(y-2)^2-5(y-2), you get y^2-9y+14. When you plug in y-2 into h(y-2)=2(y-2)+1, you get 2y-3 The question asks g(y-2)-h(y-2), so y^2-9y+14 - 2y-3 I get y^2-11y+11, and that's wrong.

OpenStudy (anonymous):

It should be y^2-11y+17 but I don't know how

OpenStudy (anonymous):

The question asks g(y-2)-h(y-2), so y^2-9y+14 - 2y+3=y^2-11y+17 -----

OpenStudy (anonymous):

how did you get positive 3?

OpenStudy (anonymous):

Ooh, so I distribute the minus sign in g(y-2) - h(y-2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ur welcome and \[\huge Welcome \ \ 2 \ \ Open \ \ Study\]

OpenStudy (jiteshmeghwal9):

As we know that in function if we put any value in bracket then, we also have to put the value in the given expression. now it is given to us that \[g(n)=n^2-5n , h(n)=2n+1\] \[To find :-g(y-2)-h(y-2)\] \[g(y-2)=(y-2)^2-5n , h(y-2)=2(y-2)+1\] now, put this values of g(y-2) & h(y-2) as follows\[((y-2)^2-5n)-(2y-4-1)\] now, solve ths & get ur answer:)

OpenStudy (jiteshmeghwal9):

Sorry, \[g(y-2)=(y-2)^2-5(y-2)\] so, u have to solve \[((y-2)^2-5(y-2)-(2y-4-1)\]